OpenStudy (jkbo):

Please help me solve Write y = 2/3 x + 7 in standard form using integers. A.–2x + 3y = 21 B.–2x – 3y = 21 C.3x – 2y = 21 D.–2x + 3y = 7

OpenStudy (akashdeepdeb):

\[ y = 2/3 x + 7 ~~~~~ =>~~~~ y = \frac{ 2 }{ 3 }x+ \frac{ 7*3 }{ 3 }\] \[~~~~=>~~~~~~~~~~y = \frac{ 2x + 21 }{ 3 }~~~~~~~=> ~~~~~~3y = 2x + 21\] \[~~~~=>~~~~~~~~~~-2x + 3y = 21\] Understood? :)

OpenStudy (jkbo):

Thank you very much

OpenStudy (akashdeepdeb):

:)